OpenStudy (anonymous):

simplify to lowest term x^2-5x+4/x^2+4x-32 I was able to get (x+5)(x+4) is this correct

OpenStudy (cwrw238):

x^2-5x+4 = (x -4)(x - 1) x^2+ 4x - 32 = (x + 8)(x - 4) so we have